how would i solve this by hand? i know from the calcualator the answer is -3 but i'd like to learn how to do it the other way.
-2 ln e^2 + 1

Answer by ankor@dixie-net.com(12692)

-2 ln e^2 + 1
Write the log equiv of exponents: ln(e^2) = 2*ln(e)
So you have:
-2 * 2*ln(e) + 1
Which is
-4*ln(e) + 1
The nat log of e = 1, therefore
-4*1 + 1
-4 + 1 = -3